OpenStudy (anonymous):

The sum of three consecutive integers is 267. What is the largest integer?

OpenStudy (agent0smith):

Call the largest integer x. You can write three consecutive integers as (from biggest to smallest): x, x-1, x-2. They add up to 267, so\[ \Large x + x-1 + x-2 = 267\] Solve for x.

OpenStudy (wolf1728):

x + (x+1) + (x+2) = 267 3x + 3 = 267 3 x = 264 x = 88 So the highest integer = 90

OpenStudy (agent0smith):

The reason you make them x, x-1, x-2, is because if, for example, the largest integer is 100, the next is 99 (which is 100 - 1), then 98 (which is 100 -2).

OpenStudy (agent0smith):

@wolf1728 give him a chance to at least attempt it before posting the answer...

OpenStudy (wolf1728):

agent0smith Sorry for posting too quickly (I've been doing that a lot too) I'll try to calm it down just a bit.